OpenStudy (anonymous):

if sin(theta) = 2/3 find csc(theta)

OpenStudy (anonymous):

\[\sin \Theta = \frac{ 2 }{ 3 } find \csc \Theta \]

OpenStudy (anonymous):

given cos theta = 2/3 ans sin theta > 0 find: sin theta = (sqrt(5))/2 tan theta = (sqrt(5))/6 cot theta = sec theta = csc theta = 3/(sqrt(5))